# Merging Facts, Crafting Fallacies: Evaluating the Contradictory Nature of Aggregated Factual Claims in Long-Form Generations

### 08 Feb 2024 v1

@@ Abstract
Long-form generations from large language models (LLMs) contains a mix of factual and non-factual claims, making evaluating factuality difficult. To evaluate factual precision of long-form generations in a more fine-grained way, prior works propose to decompose long-form generations into multiple verifiable facts and verify those facts independently. The factuality of the generation is the proportion of verifiable facts among all the facts. Such methods assume that combining factual claims forms a factual paragraph. This paper shows that the assumption can be violated due to entity ambiguity. We show that LLMs can generate paragraphs that contain verifiable facts, but the facts are combined to form a non-factual paragraph due to entity ambiguity. We further reveal that existing factual precision metrics, including FActScore and citation recall, cannot properly evaluate the factuality of these non-factual paragraphs. To address this, we introduce an enhanced metric, D-FActScore, specifically designed for content with ambiguous entities. We evaluate the D-FActScores of people biographies generated with retrieval-augmented generation (RAG). We show that D-FActScore can better assess the factuality of paragraphs with entity ambiguity than FActScore. We also find that four widely used open-source LLMs tend to mix information of distinct entities to form non-factual paragraphs.
@@ 1 Introduction
Large language models (LLMs) can generate fluent and natural texts, making LLMs prevalent in everyday usage (Brown et al., 2020; OpenAI, 2022, 2023). However, LLM-generated texts may not always have a high factual precision (Nakano et al., 2021; Rae et al., 2021). Factual precision measures whether the information conveyed in the text is factually accurate. [^1] In this paper, we only focus on factual precision and do not consider factual recall, i.e., how well the generation covers the information. This paper will use factuality to refer to factual precision. As long-form generations can contain a mix of factual and non-factual information, recent works propose to evaluate the factuality of long-form generation in a more fine-grained way that considers the factuality of each claim in the generation. Precisely, these methods decompose the generation into several claims and verify each claim independently. The factuality of the whole generation is the percentage of verifiable claims. Widely used evaluation metrics like FActScore (Min et al., 2023a) and citation recall (Liu et al., 2023; Gao et al., 2023) fall into this type.
Figure 1: A paragraph generated by Llama-13b-chat when prompted to generate a biography for Dick Hanley. We provide top-5 documents retrieved from Wikipedia for RAG. While the generated paragraph is non-factual, each fact is supported by Wikipedia.
@@
The correctness of the factuality metrics mentioned above relies on the assumption that "as long as each claim is factual, the combination of those claims is factual". This paper shows that this assumption can be violated due to entity ambiguity in the generation. Figure 1 shows such a case, where a paragraph mixes the information of two entities with the same name. In this example, each claim in the generation is factual because it can be supported Wikipedia. As a result, FActScore of the above misleading paragraph is 100% by definition.
@@
We emphasize the significance of this evaluation problem by showing that LLMs can easily generate this kind of non-factual paragraph. We prompt LLMs to generate a biography for a target name using retrieval-augmented generation (Guu et al., 2020; Lewis et al., 2020), where the target name is shared by multiple entities, and the retrieved documents may also include the Wikipedia pages of multiple entities. We find that Llama-13b/70b-chat (Touvron et al., 2023), Tulu-13b-dpo (Ivison et al., 2023), and Vicuna-7b (Chiang et al., 2023) tend to mix the information of multiple individuals in the same biography, and a reader without prior knowledge cannot tell that the information corresponds to different people. Non-factual as these paragraphs are, existing factuality metrics cannot correctly assess the factuality in this case.
We modify FActScore into Disambig-FActScore (D-FActScore), which handles entity ambiguity better. Unlike FActScore, which verifies each fact independently, D-FActScore splits the facts in a paragraph into groups based on disambiguation information in the paragraph and considers the factuality of the facts within the group altogether.
@@
    We show that open-source LLMs cannot yet properly handle entity ambiguity in the retrieved passages to generate a factual paragraph, yielding much lower D-FActScores compared to ChatGPT.
@@ 2 LLMs Combine Factual Claims into Non-factual Paragraph
This section shows how to prompt LLMs to generate people biographies full of factual claims but combined in a non-factual way due to entity ambiguity. We choose to generate people biographies for the following reasons:
  • The factuality of biographies can be easily verified due to the broad coverage of Wikipedia.
  • There are extensive prior works that use biography generation to evaluate the factuality of LLMs (Min et al., 2023a; Asai et al., 2023).
  • Given that name-based queries are common in online searches, it is likely that LLM users will directly request information about specific entities from LLMs.
@@ 2.1 Collecting Names with Ambiguity
To generate biographies with mixed information of multiple entities with the same name, we collect names with ambiguity from Wikipedia disambiguation pages using the Wikipedia API. A disambiguation page lists all the Wikipedia pages of the entities with the same name. We randomly select 500 ambiguous names and call this collection AmbigBio in our paper and use target name to refer to a name in AmbigBio. A target name may correspond to multiple entities in Wikipedia.
@@ 2.2 Prompting LLMs to Generate Biographies
We prompt LLMs to generate biographies with retrieval-augmented generation.
Retrieval.   Given a target name in AmbigBio, we retrieve the top-5 passages from Wikipedia using GTR (Ni et al., 2022). The knowledge source for retrieval is the 2018-12-20 Wikipedia snapshot split into passages of 100 words (Karpukhin et al., 2020). The query used for retrieval is "Tell me a bio of <entity>", where <entity> is a target name. Some names in AmbigBio contain the string "(disambiguation)", indicating that it is a disambiguation page. The "(disambiguation)" is removed when creating the query.
@@
Name of the person: <entity>
@@
Generation.   We prompt LLMs with the top-5 retrieved passages to generate biographies for a target name with citations (Nakano et al., 2021; Gao et al., 2023). The LLM must cite the retrieved Wikipedia passages in the generation for attribution. The reason to generate biographies with citations is to evaluate the attribution of the generated content. In Appendix D.1, we will show that even perfect citation attribution can still be non-factual. We prompt the LLMs by Vanilla prompting used in Gao et al. (2023) due to its superior performance in generating text with citations. A simplified prompt is shown in Table 1. The title for each retrieved passage is included in the prompt. The titles in Wikipedia sometimes contain parenthesis for disambiguation, e.g., Dick Hanley (Swimmer) in Figure 1. We do not remove the words in the parenthesis when prompting the LLM, allowing the LLM to use this information for disambiguation.
@@
(1) With name ambiguity: The person names in the demonstrations are ambiguous names in Wikipedia, and the retrieved passages include Wikipedia pages of different entities with the same name. The demonstration biography only contains the information of one of the entities with the target name.
(2) Without name ambiguity: The name is unambiguous, in the sense that there is no disambiguation page for that name. The retrieved results contain the passages from that entity’s Wikipedia page and possibly some unrelated passages. The demonstration answer is the biography of that entity.
@@ 2.3 Large Language Models
We use five LLMs in our paper: Llama-13b-chat, Llama-70b-chat (Touvron et al., 2023), Vicuna-7b (Chiang et al., 2023), Tulu-v2-13b-dpo (Ivison et al., 2023), and ChatGPT (gpt-3.5-turbo-0301(OpenAI, 2022). These LLMs differ in size, alignment methods, and whether they are open-sourced or not.
2.4 Qualitivate Analysis of LLM Outputs
We categorize the generated paragraphs based on the number of distinct entities and the number of disambiguable biographies, defined as follows. The below definitions and categorization are better understood with the examples in Table 2. We also formally define some frequently used term.
Definition: (Named) Entity. A named entity is an object in real-world that can be denoted with a proper noun. In our paper, an entity is a real-world human with a corresponding Wikipedia page.
Definition: (Entity) Mention. An entity mention is a specific instance when a named entity is referenced or mentioned within text. For the same entity, it may be mentioned in different ways.
Definition: Number of distinct entities. This is the count of distinct entities with the same name as the target name in an LLM-generated paragraph. This can be calculated by performing entity linking to mentions with the target name in the paragraph and counting how many distinct entities in Wikipedia are linked.
Definition: Number of distinguishable biographies. A paragraph might contain information about multiple entities, but its wording could mislead readers into believing it is about only one individual, making it seem like a single biography (as shown in Figure 1). The number of distinguishable biographies is given by the following operational definition: How many individuals with the target name willone identify in the paragraph without relying on prior or external knowledge? Note that a paragraph can contain more than one biography.
We categorize the LLM outputs into four types:
Type LLM generation
@@ 2.4 Qualitivate Analysis of LLM Outputs
Table 2: Examples of three different types of LLM outputs. [Bio $i$] denotes the start of the $i$-th biography. [Ent $i$] denotes the subsequent information is about the $i$-th entity in Wikipedia that has the target name. We remove the citations ([1][2][3]) from the outputs here.
One biography and one entity
The generated output only contains the information of one entity. In other words, all the information in the generated output points to the same Wikipedia entity whose name is the target name.
One biography and $N$ entities
The generated output mixes the information of different entities with the target name but does not provide sufficient disambiguation information. A typical reader without prior knowledge will consider the whole paragraph a single biography of one person.
$N$ biographies and $N$ entities
The generated output contains information about multiple entities, and the paragraph provides enough context for readers without prior knowledge to disambiguate different entities in the paragraph. Each biography is about one of the entities with the target name.
No biography and no entity
This is the case when the LLM abstains from answering, which can happen when the retrieved passages do not contain the information of individuals with the target name or the LLM finds multiple entities with the same name and asks for disambiguation.
Ideally, the number of distinguishable biographies should match the number of distinct entities. When these two numbers agree, the LLM-generated paragraph provides enough disambiguation information in the paragraph and is likely factual. Additionally, when multiple valid entities exist in the retrieved documents, LLMs should generate a paragraph that includes all the entities. This informs the users about the name’s ambiguity and improves the chances that the content generated by the LLMs covers the entity who the user is asking about.
@@ 3 Existing Factual Precision Metrics
We discuss why some factual precision metrics cannot properly assess the factuality of paragraphs with entity ambiguity, as shown in Section 2.
@@
The key idea of FActScore is to decompose a long-form generation $y$ into a list of atomic facts $\mathcal{A}_{y}$ short sentences that convey one piece of information. FActScore of a long-form generation $y$ is defined by
@@
Min et al. (2023a) show that FActScore can be obtained using an automatic pipeline: they instruct text-davinci-003 to decompose a long-form generation into atomic facts and use another $\text{LM}_{\text{EVAL}}$ to verify each fact. They propose three different types of $\text{LM}_{\text{EVAL}}$ to verify an atomic fact: (1) No-context LM: prompt an $\text{LM}_{\text{EVAL}}$ using ’<atomic fact> True or False?’ (2) Retrieve $\to$ LM: retrieve $k$ passages from Wikipedia, construct the prompt by concatenating the retrieved passages, the atomic fact, and ’True or False?’, and prompt $\text{LM}_{\text{EVAL}}$ to answer. (3) Nonparametric Probability (NP): this is calculated by the average token likelihood of the atomic fact using a nonparametric masked LM (Min et al., 2023b). The atomic fact is considered True if the average token likelihood is higher than a pre-defined threshold. Min et al. (2023a) show that ensembling NP and Retrieve $\to$ LM can approximate the FActScore obtained by human annotation.
Shortcoming of FActScore
FActScore is not designed to handle entity ambiguity because FActScore considers the factuality of each atomic fact independently without considering the whole paragraph. We explain this using the example in Figure 1. In the Retrieve $\to$ LM method, each atomic fact about Dick Hanley in the paragraph is verifiable by Wikipedia, leading to a 100% FActScore. However, some facts are supported by Dick Hanley (American Football), and others are supported by Dick Hanley (Swimmer), indicating that there are two entities in the paragraph. But this cannot be easily seen from the paragraph itself, and readers without any prior knowledge may believe that there is one person named Dick Hanley who is an American footballer and a swimmer. Since this paragraph does not allow readers to understand the information in a factual way, the paragraph should not be considered factual.
No-context LM and NP methods also cannot properly handle entity ambiguity. For example, the fact "Dick Hanley passed away on December 16, 1970" is factual if this fact is extracted from the biography of Dick Hanley (American Football). Contrarily, if the fact is from the biography of Dick Hanley (Swimmer), the fact is non-factual. However, for the same fact, no-context LM and NP methods always yield the same result (neglecting the randomness during prediction), disregarding the context of the fact. This makes them unable to distinguish entities with the same name.
@@ 4 Disambig-FActScore (D-FActScore)
We refine FActScore into Disambig-FActScore (D-FActScore) to better address ambiguity in factuality evaluation. The definition of D-FActScore is presented in Section 4.1. Section 4.2 outlines the human annotation process for D-FActScore, with the outcomes of human evaluations detailed in Section 4.3.

### 23 Feb 2024 v2

@@ Abstract
Long-form generations from large language models (LLMs) contain a mix of factual and non-factual claims, making evaluating factuality difficult. To evaluate factual precision of long-form generations in a more fine-grained way, prior works propose to decompose long-form generations into multiple verifiable facts and verify those facts independently. The factuality of the generation is the proportion of verifiable facts among all the facts. Such methods assume that combining factual claims forms a factual paragraph. This paper shows that the assumption can be violated. We show that LLMs can generate paragraphs that contain verifiable facts, but the facts are combined to form a non-factual paragraph due to entity ambiguity. We further reveal that existing factual precision metrics, including FActScore and citation recall, cannot properly evaluate the factuality of these non-factual paragraphs. To address this, we introduce an enhanced metric, D-FActScore, specifically designed for content with ambiguous entities. We evaluate the D-FActScores of people biographies generated by retrieval-augmented LLMs. We show that D-FActScore can better assess the factuality of paragraphs with entity ambiguity than FActScore. We also find that four widely used open-source LLMs tend to mix information of distinct entities to form non-factual paragraphs.
@@ 1 Introduction
LLMs can generate high-quality texts, making LLMs prevalent in everyday usage (OpenAI, 2022, 2023). However, LLM’s generation may not always have a high factual precision (Nakano et al., 2021; Rae et al., 2021). Factual precision measures whether the information conveyed in the text is factually accurate. [^1] We only focus on factual precision and do not consider factual recall, i.e., how well the generation covers the information. This paper will use factuality to refer to factual precision. As long-form generations can contain a mix of factual and non-factual claims, recent works propose to evaluate the factuality of long-form generation in a more fine-grained way that considers the factuality of each claim in the generation. Precisely, these methods decompose the generation into several claims (called atomic facts in Min et al. (2023a)) and then verify each claim independently. The factuality of the whole generation is the percentage of verifiable claims. Widely used evaluation metrics like FActScore (Min et al., 2023a) and citation recall (Liu et al., 2023; Gao et al., 2023) fall into this type.
Figure 1: Output of an LLM when prompted to generate a biography for Dick Hanley. While the paragraph is misleading and non-factual, all the facts in the paragraph can be supported by the Wikipedia of Dick Hanley (Swimmer)  or Dick Hanley (American FootBall) , yielding 100% FActScore. D-FActScore groups atomic facts that appear to refer to the same individual based on the paragraph (Figure 1(c)), finds an entity that best matches that individual from the knowledge source (Figure 1(d)), and only uses the information of that entity to verify the facts in that group .
@@
The correctness of the factuality metrics mentioned above relies on the assumption that "as long as each claim is factual, the combination of those claims is factual". This paper shows that this assumption can be violated due to entity ambiguity in the generation. Figure 1 shows such a case, where a paragraph mixes the information of two entities with the same name, but readers without prior or external knowledge will think all the information refers to the same individual. Since the paragraph is not written in a way that allows readers to understand the content factually, the paragraph is non-factual. However, since each claim in the generation is supported by Wikipedia, the FActScore of the above misleading paragraph is 100% by definition.
@@
We emphasize the significance of this evaluation problem by showing that LLMs can easily generate this kind of non-factual paragraph. We prompt LLMs to generate a biography for a target name using retrieval-augmented generation (Guu et al., 2020; Lewis et al., 2020), where the target name is shared by multiple entities, and the retrieved documents may also include the Wikipedia pages of multiple entities. We find that Llama-13b/70b-chat (Touvron et al., 2023), Tulu-13b-dpo (Ivison et al., 2023), and Vicuna-7b (Chiang et al., 2023) tend to mix the information of multiple entities in the same biography, and a reader without prior knowledge cannot tell that the information corresponds to different entities. Non-factual as these paragraphs are, existing factuality metrics cannot correctly assess the factuality in this case.
To solve this evaluation problem, we modify FActScore into Disambig-FActScore (D-FActScore), which handles entity ambiguity better. Unlike FActScore, which verifies each atomic fact in the paragraph independently, D-FActScore splits the facts in a paragraph into groups. If the narration of the paragraph can make a reader without prior knowledge think that the two facts are about the same individual/object, they belong to the same group. Next, D-FActScore uses entity linking to find an entity in the knowledge source that best matches a group of facts and use this entity to verify the facts in the same group.
@@
    We show that four open-source LLMs cannot properly handle entity ambiguity in the retrieved passages to generate a factual paragraph, yielding much lower D-FActScores compared to ChatGPT.
@@ 2 LLMs Combine Factual Claims into Non-factual Paragraph
This section shows how to prompt LLMs to generate people biographies full of factual claims but combined in a non-factual way due to entity ambiguity. We choose to generate people biographies for the following reasons: (1) There are extensive prior works that use biography generation to evaluate the factuality of LLMs (Min et al., 2023a; Asai et al., 2023). (2) Given that name-based queries are common in online searches, it is likely that LLM users will directly request information about specific entities from LLMs.
@@ 2.1 Collecting Names with Ambiguity
To generate biographies with mixed information of multiple entities with the same name, we collect names with ambiguity from Wikipedia disambiguation pages using the Wikipedia API. A disambiguation page lists all the Wikipedia pages of the entities with the same name. We randomly select 500 ambiguous names and call this collection AmbigBio in our paper. We use target name to refer to a name in AmbigBio. A target name corresponds to multiple entities in Wikipedia.
@@ 2.2 Prompting LLMs to Generate Biographies
We prompt LLMs to generate biographies with retrieval-augmented generation (RAG).
Retrieval.   Given a target name in AmbigBio, we retrieve the top-5 passages related to the name from Wikipedia using GTR (Ni et al., 2022). The knowledge source for retrieval is the 2018-12-20 Wikipedia snapshot split into passages of 100 words (Karpukhin et al., 2020). The query used for retrieval is "Tell me a bio of <name>". Some names in AmbigBio contain the string "(disambiguation)", indicating that it is a disambiguation page. The "(disambiguation)" is removed when creating the query for retrieval.
@@
Name of the person: <name>
@@
Generation.   We prompt LLMs with the top-5 retrieved passages to generate biographies for a target name with citations (Nakano et al., 2021; Gao et al., 2023). The LLM can only use the retrieved documents and must cite the retrieved passages in the output for attribution. The reason to generate biographies with citations is to evaluate the attribution of the generated content. In Appendix D.1, we will show that even perfect citation attribution can still be non-factual. We prompt the LLMs by Vanilla prompt (shown in Table 1) used in Gao et al. (2023) due to its superior performance in generating text with citations. The title for each retrieved passage is included in the prompt. The titles in Wikipedia sometimes contain parenthesis for disambiguation, e.g., Dick Hanley (Swimmer) in Figure 1. We do not remove the words in the parenthesis when prompting the LLM to generate the biography, allowing the LLM to use this information for disambiguation.
@@
(1) With name ambiguity: The names in the demonstrations are ambiguous names in Wikipedia; the retrieved passages include Wikipedia pages of different entities with the same name. The <answer> only contains the information of one of the entities with that name.
(2) Without name ambiguity: The name in the demonstration is unambiguous (there is no disambiguation page for that name in Wikipedia). The retrieved results contain the passages from that entity’s Wikipedia page and possibly some unrelated passages. The <answer> is the bio of that entity.
@@ 2.3 Large Language Models
We use five LLMs with different sizes and alignment methods in our paper: Llama-13b-chat, Llama-70b-chat (Touvron et al., 2023), Vicuna-7b (Chiang et al., 2023), Tulu-v2-13b-dpo (Ivison et al., 2023), and ChatGPT (gpt-3.5-turbo-0301(OpenAI, 2022).
2.4 Categorizing LLM-Generated Paragraphs
We categorize the generated paragraphs based on the number of distinct entities and the number of disambiguable biographies, defined as follows. The below definitions and categorization are better understood with the examples in Table 2.
Definition: (Named) Entity. A named entity is an object in real world that can be denoted with a proper noun. In our paper, an entity is a real-world human with a corresponding Wikipedia page.
Definition: (Entity) Mention. An entity mention is a specific instance when a named entity is referenced or mentioned within text. An entity may be mentioned in different ways.
Definition: Number of distinct entities in a paragraph. A paragraph can be decomposed into atomic facts following Min et al. (2023a). Since the LLM is instructed to use only the retrieved documents to compose a paragraph, each atomic fact in the paragraph should originate from a retrieved passage, which is the Wikipedia page of an entity. We attribute each atomic fact in the paragraph to a Wikipedia entity and collect the entities into a set; the number of distinct entities is the number of unique elements in the set.
Definition: Number of distinguishable biographies in a paragraph. A paragraph might contain information about multiple entities, but its wording could mislead readers into believing it is about only one individual, making it seem like a single bio (as shown in Figure 1). The number of distinguishable bios is given by the following operational definition: How many individuals with the target name will one identify in the paragraph without relying on prior or external knowledge? Note that a paragraph can contain more than one bio.
We categorize the LLM outputs into three types:
Type LLM-generated paragraph
@@ 2.4 Categorizing LLM-Generated Paragraphs
Table 2: Examples of three different types of LLM-generated paragraphs. [Bio $i$] denotes the start of the $i$-th distinguishable biography. [Ent $i$] denotes the subsequent information is about the $i$-th entity in Wikipedia that has the target name. We remove the citations ([1][2][3]) from the LLM-generated paragraphs here.
(1) One biography and one entity: The generated output only contains the information of one entity. In other words, all the information in the generated output points to the same Wikipedia entity whose name is the target name.
(2) One biography and $N$ entities: The generated output mixes the information of different entities with the target name but does not provide sufficient disambiguation information. A typical reader without prior knowledge will consider the whole paragraph a single bio of one individual.
(3) $N$ biographies and $N$ entities: The paragraph contains information about multiple entities and provides enough context for readers without prior knowledge to disambiguate different entities in the paragraph. Each biography describes one of the entities with the target name.
Ideally, the number of distinguishable bios should match the number of distinct entities. When these two numbers agree, the LLM-generated paragraph provides enough disambiguation information in the paragraph and is likely factual.
@@ 3 Existing Factual Precision Metrics
We discuss why some factual precision metrics cannot properly assess the factuality of paragraphs with entity ambiguity, as shown in Figure 1.
@@
The key idea of FActScore is to decompose a long-form generation $y$ into a list of atomic facts $\mathcal{A}_{y}$ short sentences that convey one piece of information. FActScore of a paragraph $y$ is defined by:
@@
Min et al. (2023a) show that FActScore can be obtained using an automatic pipeline: they instruct GPT3.5 (text-davinci-003) to decompose a long-form generation into atomic facts and use another $\text{LM}_{\text{EVAL}}$ to verify each fact. They propose three different types of $\text{LM}_{\text{EVAL}}$ to verify an atomic fact: (1) No-context LM: prompt an $\text{LM}_{\text{EVAL}}$ using ’<atomic fact> True or False?’ (2) Retrieve $\to$ LM: retrieve $k$ passages from Wikipedia, construct the prompt by concatenating the retrieved passages, the atomic fact, and ’True or False?’, and prompt $\text{LM}_{\text{EVAL}}$ to answer. (3) Nonparametric Probability (NP): this is calculated by the average token likelihood of the atomic fact using a nonparametric masked LM (Min et al., 2023b). The atomic fact is considered True if the average token likelihood exceeds a pre-defined threshold. Min et al. (2023a) show that ensembling NP and Retrieve $\to$ LM can approximate the FActScore obtained by human annotation.
Shortcoming of FActScore.   FActScore is not designed to handle entity ambiguity because FActScore considers the factuality of each atomic fact independently without considering the whole paragraph. We explain this using the example in Figure 1. In the Retrieve $\to$ LM method, each atomic fact about Dick Hanley in the paragraph is verifiable by Wikipedia, leading to a 100% FActScore. However, some facts are supported by Dick Hanley (American Football), and others are supported by Dick Hanley (Swimmer), indicating that there are two entities in the paragraph. But this cannot be easily seen from the paragraph itself, and readers without any prior knowledge may believe that Dick Hanley is a swimmer who died in 1970, which is incorrect. Since this paragraph does not allow readers to understand the information in a factual way, the paragraph should not be considered factual.
No-context LM and NP methods cannot properly handle entity ambiguity. For example, consider the fact "Dick Hanley passed away on December 16, 1970". This fact is factual if it is extracted from the biography of Dick Hanley (American Football). Contrarily, if the fact is from the biography of Dick Hanley (Swimmer), the fact is non-factual. However, for the above atomic fact, no-context LM and NP methods always yield the same result and cannot consider whose biography the fact is extracted from. This makes them unable to distinguish entities with the same name.
@@ 4 Disambig-FActScore (D-FActScore)
We refine FActScore into D-FActScore to better address entity ambiguity in factuality evaluation. The definition of D-FActScore is presented in Section 4.1. Section 4.2 outlines the human annotation process for D-FActScore, with the outcomes of human evaluations detailed in Section 4.3.